# Do Determinants of EV Purchase Intent vary across the Spectrum? Evidence from Bayesian Analysis of US Survey Data

### v2

@@ 3.2 The Ordinal Quantile Model
The parameter estimates from the ordinal probit model provide information on the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates. However, researchers are often more interested in examining specific quantiles of the dependent variable, particularly the lower and upper extremes of the distribution, to answer policy-relevant questions. For instance, having confidence that public charging infrastructure will continue to expand—one of our key covariates—may have less effect on individuals with low interest in EVs. However, among those who are moderately interested, increased confidence in future infrastructure may strengthen intent, reassuring them that charging access will not be a barrier and moving them from “somewhat likely” to “very likely” buyers. To capture how the influence of infrastructure confidence and other factors varies across levels of purchase intent, we use ordinal quantile regression.
The ordinal quantile regression model, proposed by Rahman (2016), can be written in terms of a latent variable $z_{i}$ as follows:
@@
where the regression coefficient $\beta_{p}$ and cut-point vector $\gamma_{p}$ are quantile dependent as indicated by the subscript $p$, and $\varepsilon_{i}\overset{iid}{\sim}AL(0,1,p)$ for $i=1,2,\ldots,n$; where AL denotes an asymmetric Laplace distribution (Yu and Zhang, 2005). The AL distribution is essential to create a working likelihood function, and yields the following expression,
$\begin{split}f(y|\beta_{p},\gamma_{p})&=\prod_{i=1}^{n}\prod_{j=1}^{J}P(y_{i}=j|\beta_{p},\gamma_{p})^{I(y_{i}=j)}\\ &=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg[F_{AL}(\gamma_{p,j}-x^{\prime}_{i}\beta_{p})-F_{AL}(\gamma_{p,j-1}-x^{\prime}_{i}\beta_{p})\bigg]^{I(y_{i}=j)},\end{split}$ (7)
@@
At this stage, three key issues need attention. First, the ordering constraint of cut-points during sampling is addressed using the transformation, $\delta_{p,j}=\ln(\gamma_{p,j}-\gamma_{p,j-1})$ for $2\leq j\leq J-1$. Second, we set $\gamma_{p,1}=0$ to anchor the location and variance is automatically fixed for any quantile to meet the scale restriction. Third, the AL distribution is not convenient for MCMC sampling as it does not yield tractable conditional posteriors. So, following Kozumi and Kobayashi (2011), we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$. The variable $w_{i}$ follows an exponential distribution, $w_{i}\sim\mathcal{E}(1)$, and is independent of $u_{i}$, which follows a normal distribution, $u_{i}\sim N(0,1)$. Based on this formulation, $z_{i}|\beta_{p},w_{i}\sim N(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\ldots,n$; enabling us to leverage the properties of normal distribution for an efficient MCMC algorithm.
@@
$\begin{split}\pi(z,\beta_{p},\delta_{p},w|y)&\propto f(y|z,\beta_{p},\delta_{p},w)\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\Big\{\prod_{i=1}^{n}f(y_{i}|z_{i},\delta_{p})\Big\}\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\prod_{i=1}^{n}\bigg\{\prod_{j=1}^{J}1\{\gamma_{p,j-1}<z_{i}<\gamma_{p,j}\}\;N(z_{i}|x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})\;\mathcal{E}(w_{i}|1)\bigg\}\\ &\qquad\times\;N(\beta_{p}|\beta_{p0},B_{p0})\;N(\delta_{p}|\delta_{p0},D_{p0}).\end{split}$ (8)
In the first line, we use the factorization $f(y,z|\beta_{p},\delta_{p},w)=f(y|z,\beta_{p},\delta_{p},w)\times f(z|\beta_{p},w)$ and note that $z$ is independent of $\delta_{p}$. In the second line, we exploit the independence of $y_{i}$ from $(\beta_{p},w)$ given $(z_{i},\delta_{p})$, which follows from the second line of equation (6), where $y_{i}$ given $(z_{i},\delta_{p})$ is determined with probability 1. In the third and fourth lines, we specify the conditional distribution of the latent variable $z$ and the prior distribution on the parameters $(\beta_{p},\delta_{p})$.
@@
    otherwise repeat the value $\delta_{p}$. The variance of $s$ may be tuned for appropriate step size and acceptance rate.
@@
The conditional posterior distributions are derived from the complete data posterior (i.e., equation 8) and the parameters are sampled as per Algorithm 2. The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman (2023).
After estimating the ordinal quantile models, we use the method of composition to compute the covariate effects, such as in (Rahman and Vossmeyer, 2019) and Bresson et al. (2021). Assuming a setup similar to Section 3.1, our goal is to draw from the distribution $\{\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})\}$. So, we randomly select an individual, extract their covariates, draw a value $(\beta_{p},\delta_{p})$ from their posterior distributions, and finally evaluate $\{\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta_{p},\delta_{p})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta_{p},\delta_{p})\}$, where now for the quantile model: $\Pr(y_{i}=j|x_{i}^{q},\beta_{p},\delta_{p})=F_{AL}(\gamma_{p,j}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})-F_{AL}(\gamma_{p,j-1}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})$, for $q=b,a$ and $1\leq j\leq J$. This process is repeated for all individuals and MCMC draws, and the average covariate effect for outcome $j$ at quantile $p$ is calculated as the mean of difference in pointwise probabilities.
@@ 4.1 Results from Ordinal Probit Models
We begin by presenting the results from ordinal probit models estimated separately for each survey. Table 2 presents the posterior means and standard deviations of the parameters based on 10,000 MCMC iterations after a burn-in of 2,500 iterations. We utilize the priors $\beta\sim N_{k}(0,I)$ and $\delta\sim N_{J-2}(0,0.25\times I)$, where $I$ denotes an identity covariance matrix. These priors are relatively diffuse, allowing the data to primarily drive the posterior inference. Table 2 also reports the logarithm of marginal likelihood, which is the most accepted quantity for Bayesian model comparison (Chib and Jeliazkov, 2001; Jeliazkov et al., 2008). The marginal likelihood (also known as model evidence) is the probability of observed data, obtained by integrating the likelihood function with respect to prior distribution. It accounts for both model fit and complexity, with higher values indicating a better model fit. The use of marginal likelihood contrasts with those commonly used in the classical framework, such as the pseudo-$R^{2}$, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) (Lin and Tan, 2017; Kumar et al., 2025).
@@
The combined results from the three surveys reveal consistent patterns in factors associated with EV purchase intent. As shown in Table 3, the covariate effects from the 2021 survey show that respondents who have read or heard extensively about EVs are 10.2 (2.2) percentage point more likely to express strong interest (some interest) in purchasing EVs, compared to those with limited exposure. These gains in probability comes from a reduction of 8.5 and 3.9 percentage points from the first two opinion categories. Our findings complements Zhao et al. (2022) and Scherrer (2023), who demonstrate that greater exposure to EV information through peers or media positively influences attitudes toward EVs, and with Krause et al. (2013), who underscore the role of accurate information in shaping perceptions.
Going forward, we discuss the covariate effects only on the first and last categories as they show the largest changes in probability. For the 2021 survey, the belief that EVs are environmentally better than gas-powered vehicles is associated with a 15.5 percentage point increase in strong interest and a 19.5 percentage point decrease in no interest, underscoring the importance of environmental perceptions in shaping purchase intent. This finding complements Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025) who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs. Here the effect reflects perceptions of EVs’ benefits over gasoline vehicles rather than general pro-environmental attitudes [^7] Relatedly, Buhmann and Criado (2023) find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
@@
Confidence in charging infrastructure, assessed in the 2023 survey, is one of the strongest predictors of EV interest. Respondents very or extremely confident that the US will build the necessary infrastructure are 19.1 (18.8) percentage points more likely to express strong (no) interest in purchasing EVs, compared to those who report lower confidence. This finding complements research, such as Carley et al. (2013) and Li et al. (2017), which highlight the importance of existing infrastructure.
@@
Across all surveys, previous EV ownership is consistently linked with a higher purchase interest. Respondents who have owned an EV or hybrid are 19.0–23.1 percentage points more likely to report strong interest in buying an EV and 13.2–18.1 percentage points less likely to express no interest than those without prior ownership. This pattern, also noted by Carley et al. (2013) and Buhmann and Criado (2023), suggests that firsthand experience reduces uncertainty and fosters positive perceptions of EVs. However, selection effects may also be at play, as these individuals are likely early adopters with a strong pre-existing interest in EVs.
@@
Respondents under age 50 are 2.8–3.7 percentage points more likely to express strong interest and 3.3–4.1 percentage points less likely to report no interest in purchasing EVs than older respondents. Our finding aligns with Zhao et al. (2022), Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025). For instance, Zhao et al. (2022) finds that EV buyers in Shanghai are mostly aged 31-40.
@@
Results from the 2022 and 2023 surveys also show that female respondents are are 2.3–2.4 percentage points less likely to express strong interest, and 2.7–3.0 percentage points more likely to report no interest, than males. This finding is consistent with Carley et al. (2013), Krause et al. (2013), Scherrer (2023), Buhmann and Criado (2023), and Mamkhezri (2025), which also find that men show stronger preference for EVs.
@@
Higher education is also linked to greater EV interest. Respondents with a graduate degree are 8.6–10.0 percentage points more likely to report strong interest and 8.1–12.6 percentage points less likely to express no interest than those with a high school diploma or less. Similar, though smaller, effects are observed for respondents with some college education. Such a relationship is similarly documented in Carley et al. (2013), Krause et al. (2013), Buhmann and Criado (2023), and Mamkhezri (2025).
@@
Metropolitan residents are 3.6–5.3 percentage points more likely to report strong interest and 5.1–6.6 percentage points less likely to express no interest. This finding aligns with Buhmann and Criado (2023) and Mamkhezri (2025), who also report higher EV preference among urban residents. Moreover, respondents in the West are 3.6–4.4 percentage points more likely to report strong interest than those in the Northeast; those in the Midwest and South show no significant difference from Northeastern respondents. This pattern likely reflects California’s leadership in EV adoption, driven in part ZEV mandates, the Clean Vehicle Rebate Program, and a robust charging network. [^8] See https://www.eia.gov/todayinenergy/detail.php?id=61082
@@
US-born respondents are 4.4–7.8 percentage points less likely to express strong interest and 4.2–7.8 percentage points more likely to express no interest than non-US-born respondents. Moreover, Black respondents are 2.9–5.9 percentage points less likely to express strong interest and 3.5–6.1 percentage points more likely to express no interest, as compared to White respondents. Other non-white respondents show similar preferences to White respondents. These differences may reflect cultural attitudes, driving patterns, cost perceptions, or structural barriers such as limited charging access (Hsu and Fingerman, 2021).
@@
Political affiliation is a consistent and significant predictor of EV purchase intent. Democrats are 13.9–18.8 percentage points more likely to report strong interest in purchasing EVs and 12.1–18.8 percentage points less likely to report no interest than Republicans. Independents and others are about 8 percentage points more likely to express strong interest than Republicans. These findings align with Mamkhezri (2025) and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible, as documented by Sintov et al. (2020). They also reflect the stark policy differences between the two major political parties, particularly during the tenures of Presidents Biden and Trump. During his first term (2017–2021), Trump rolled back Obama-era fuel efficiency standards, slowing regulatory pressure on automakers to transition toward EVs. In his second term, the Trump administration has again taken a deregulatory approach. In contrast, the Biden administration adopted an aggressive pro-EV stance, prioritizing stricter fuel efficiency and emissions standards and allocating \$7.5 billion through the Infrastructure Investment and Jobs Act to develop a nationwide network of 500,000 EV charging stations.
@@
Income has a positive effect on EV purchase intent. Conditional on other factors, an additional \$50,000 in income changes the probability of strong interest (no interest) by 2.1–2.9 (2.2–3.6) percentage points. Our results find support in Zhao et al. (2022) and Mamkhezri (2025), who report greater support for EVs among higher-income individuals. The small covariate effect may reflect the growing availability of EVs across various market segments and the subsidies for low-income consumers, which reduce affordability barriers compared to earlier market years.

### v3

@@ 3.2 The Ordinal Quantile Model
The parameter estimates from the ordinal probit model characterize the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates. However, policy-relevant questions often require examining how covariates affect specific parts of the outcome distribution, particularly the lower and upper extremes. For instance, having confidence that public charging infrastructure will continue to expand—one of our key covariates—may have less effect on individuals with low interest in EVs. However, among those who are moderately interested, increased confidence in future infrastructure may strengthen intent, reassuring them that charging access will not be a barrier and moving them from “somewhat likely” to “very likely” buyers. To capture such heterogeneity in covariate effects across different levels of purchase intent, we employ ordinal quantile regression.
The ordinal quantile regression model, proposed by Rahman [46], can be written in terms of a latent variable $z_{i}$ as follows:
@@
where the regression coefficient $\beta_{p}$ and cut-point vector $\gamma_{p}$ are quantile dependent as indicated by the subscript $p$, and $\varepsilon_{i}\overset{iid}{\sim}AL(0,1,p)$ for $i=1,2,\ldots,n$; where AL denotes an asymmetric Laplace distribution [57]. The AL distribution is essential to create a working likelihood function, and yields the following expression,
$\begin{split}f(y|\beta_{p},\gamma_{p})&=\prod_{i=1}^{n}\prod_{j=1}^{J}\bigg[F_{AL}(\gamma_{p,j}-x^{\prime}_{i}\beta_{p})-F_{AL}(\gamma_{p,j-1}-x^{\prime}_{i}\beta_{p})\bigg]^{I(y_{i}=j)},\end{split}$ (7)
@@
At this stage, three key issues need attention. First, the ordering constraint of cut-points during sampling is addressed using the transformation, $\delta_{p,j}=\ln(\gamma_{p,j}-\gamma_{p,j-1})$ for $2\leq j\leq J-1$. Second, we set $\gamma_{p,1}=0$ to anchor the location and variance is automatically fixed for any quantile to meet the scale restriction. Third, the AL distribution is not convenient for MCMC sampling as it does not yield tractable conditional posteriors. So, following Kozumi and Kobayashi [32], we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$. The variable $w_{i}$ follows an exponential distribution, $w_{i}\sim\mathcal{E}(1)$, and is independent of $u_{i}$, which follows a normal distribution, $u_{i}\sim N(0,1)$. Based on this formulation, $z_{i}|\beta_{p},w_{i}\sim N(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\ldots,n$; enabling us to leverage the properties of normal distribution for an efficient MCMC algorithm.
@@
$\begin{split}&\pi(z,\beta_{p},\delta_{p},w|y)\propto f(y|z,\beta_{p},\delta_{p},w)\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\Big\{\prod_{i=1}^{n}f(y_{i}|z_{i},\delta_{p})\Big\}\;f(z|\beta_{p},w)\;\pi(w)\;\pi(\beta_{p})\;\pi(\delta_{p}),\\ &\propto\prod_{i=1}^{n}\bigg\{\prod_{j=1}^{J}1\{\gamma_{p,j-1}<z_{i}<\gamma_{p,j}\}\;N(z_{i}|x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})\;\mathcal{E}(w_{i}|1)\bigg\}\pi(\beta_{p})\pi(\delta_{p}),\end{split}$ (8)
where $\pi(\beta_{p})$ and $\pi(\delta_{p})$ denote the normal priors. In the first line, we use the factorization $f(y,z|\beta_{p},\delta_{p},w)=f(y|z,\beta_{p},\delta_{p},w)\times f(z|\beta_{p},w)$ and note that $z$ is independent of $\delta_{p}$. In the second line, we exploit the independence of $y_{i}$ from $(\beta_{p},w)$ given $(z_{i},\delta_{p})$, which follows from the second line of equation (6), where $y_{i}$ given $(z_{i},\delta_{p})$ is determined with probability 1. In the third and fourth lines, we specify the conditional distribution of the latent variable $z$ and the prior distribution on the parameters $(\beta_{p},\delta_{p})$.
@@
    else repeat value $\delta_{p}$. The variance of $s$ may be tuned to achieve desired acceptance rate.
@@
The conditional posterior distributions are derived from the complete data posterior (i.e., equation 8) and the parameters are sampled as per Algorithm 2. The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman [41].
After estimating the ordinal quantile models, we use the method of composition to compute the covariate effects, such as in [45] and Bresson et al. [7]. Assuming a setup similar to Section 3.1, our goal is to draw from the distribution $\{\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})\}$. So, we randomly select an individual, extract their covariates, draw a value $(\beta_{p},\delta_{p})$ from their posterior distributions, and finally evaluate $\{\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta_{p},\delta_{p})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta_{p},\delta_{p})\}$, where now for the quantile model: $\Pr(y_{i}=j|x_{i}^{q},\beta_{p},\delta_{p})=F_{AL}(\gamma_{p,j}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})-F_{AL}(\gamma_{p,j-1}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})$, for $q=b,a$ and $1\leq j\leq J$. This process is repeated for all individuals and MCMC draws, and the average covariate effect for outcome $j$ at quantile $p$ is calculated as the mean of difference in pointwise probabilities.
@@ 4.1 Results from Ordinal Probit Models
We begin by presenting the results from ordinal probit models estimated separately for each survey. Table 2 presents the posterior means and standard deviations of the parameters based on 10,000 MCMC iterations after a burn-in of 2,500 iterations. We utilize the priors $\beta\sim N_{k}(0,I)$ and $\delta\sim N_{J-2}(0,0.25\times I)$, where $I$ denotes an identity covariance matrix. These priors are relatively diffuse, allowing the data to primarily drive the posterior inference. Table 2 also reports the logarithm of marginal likelihood [13, 25], which can be used for Bayesian model comparison. The marginal likelihood (or model evidence) is the probability of observed data, obtained by integrating the likelihood function with respect to prior distribution. It accounts for both model fit and complexity, with higher values indicating stronger support for a model [^6] In Table 2, we fit an ordinal probit model (say M1) to the 2021 survey data and obtain logarithm of marginal likelihood as $\ln f(y|M1)=-12415$. Suppose, we now estimate a reduced model (denoted M0) that excludes the regional indicators, for which the logarithm of marginal likeilhood is $\ln f(y|M0)=-12425$. Then the logarithm of Bayes Factor (BF) in favor of M1 relative to M0 is, $\ln(BF)_{10}=-12415-(-12425)=10$. This constitutes strong evidence in favor of the model with regional indicators. We note that marginal likelihoods (and hence Bayes factors) are only comparable across models fitted to the same dataset; consequently, comparisons across different survey years are not meaningful, as the models are conditioned on different observed data.. The use of marginal likelihood contrasts with commonly used classical measures, such as the pseudo-$R^{2}$, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) [38, 34].
@@
The combined results from the three surveys reveal consistent patterns in factors associated with EV purchase intent. As shown in Table 3, the covariate effects from the 2021 survey show that respondents who have read or heard extensively about EVs are 10.2 (2.2) percentage point more likely to express strong interest (some interest) in purchasing EVs, compared to those with limited exposure. These gains in probability comes from a reduction of 8.5 and 3.9 percentage points from the first two opinion categories. Our findings complements [58] and Scherrer [49], who demonstrate that greater exposure to EV information through peers or media positively influences attitudes toward EVs, and with Krause et al. [33], who underscore the role of accurate information in shaping perceptions.
Going forward, we discuss the covariate effects only on the first and last categories as they show the largest changes in probability. For the 2021 survey, the belief that EVs are environmentally better than gas-powered vehicles is associated with a 15.5 percentage point increase in strong interest and a 19.5 percentage point decrease in no interest, underscoring the importance of environmental perceptions in shaping purchase intent. This finding complements Carley et al. [10], Krause et al. [33], Scherrer [49], and Mamkhezri [42] who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs. Here the effect reflects perceptions of EVs’ benefits over gasoline vehicles rather than general pro-environmental attitudes [^8] Relatedly, Buhmann and Criado [8] find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
@@
Confidence in charging infrastructure, assessed in the 2023 survey, is one of the strongest predictors of EV interest. Respondents very or extremely confident that the US will build the necessary infrastructure are 19.1 (18.8) percentage points more likely to express strong (no) interest in purchasing EVs, compared to those who report lower confidence. This finding complements research, such as Carley et al. [10] and Li et al. [37], which highlight the importance of existing infrastructure.
@@
Across all surveys, previous EV ownership is consistently linked with a higher purchase interest. Respondents who have owned an EV or hybrid are 19.0–23.1 percentage points more likely to report strong interest in buying an EV and 13.2–18.1 percentage points less likely to express no interest than those without prior ownership. This pattern, also noted by Carley et al. [10] and Buhmann and Criado [8], suggests that firsthand experience reduces uncertainty and fosters positive perceptions of EVs. However, selection effects may also be at play, as these individuals are likely early adopters with a strong pre-existing interest in EVs.
@@
Respondents under age 50 are 2.8–3.7 percentage points more likely to express strong interest and 3.3–4.1 percentage points less likely to report no interest in purchasing EVs than older respondents. Our finding aligns with Zhao et al. [58], Carley et al. [10], Krause et al. [33], Scherrer [49], and Mamkhezri [42]. For instance, Zhao et al. [58] finds that EV buyers in Shanghai are mostly aged 31-40.
@@
Results from the 2022 and 2023 surveys also show that female respondents are are 2.3–2.4 percentage points less likely to express strong interest, and 2.7–3.0 percentage points more likely to report no interest, than males. This finding is consistent with Carley et al. [10], Krause et al. [33], Scherrer [49], Buhmann and Criado [8], and Mamkhezri [42], which also find that men show stronger preference for EVs.
@@
Higher education is also linked to greater EV interest. Respondents with a graduate degree are 8.6–10.0 percentage points more likely to report strong interest and 8.1–12.6 percentage points less likely to express no interest than those with a high school diploma or less. Similar, though smaller, effects are observed for respondents with some college education. Such a relationship is similarly documented in Carley et al. [10], Krause et al. [33], Buhmann and Criado [8], and Mamkhezri [42].
@@
Metropolitan residents are 3.6–5.3 percentage points more likely to report strong interest and 5.1–6.6 percentage points less likely to express no interest. This finding aligns with Buhmann and Criado [8] and Mamkhezri [42], who also report higher EV preference among urban residents. Moreover, respondents in the West are 3.6–4.4 percentage points more likely to report strong interest than those in the Northeast; those in the Midwest and South show no significant difference from Northeastern respondents. This pattern likely reflects California’s leadership in EV adoption, driven in part ZEV mandates, the Clean Vehicle Rebate Program, and a robust charging network. [^9] See https://www.eia.gov/todayinenergy/detail.php?id=61082
@@
US-born respondents are 4.4–7.8 percentage points less likely to express strong interest and 4.2–7.8 percentage points more likely to express no interest than non-US-born respondents. Moreover, Black respondents are 2.9–5.9 percentage points less likely to express strong interest and 3.5–6.1 percentage points more likely to express no interest, as compared to White respondents. Other non-white respondents show similar preferences to White respondents. These differences may reflect cultural attitudes, driving patterns, cost perceptions, or structural barriers such as limited charging access [21].
@@
Political affiliation is a consistent and significant predictor of EV purchase intent. Relative to Republicans, Democrats are 13.9–15.6 percentage points more likely to report strong interest in purchasing an EV and 12.1–18.8 percentage points less likely to report no interest, while Independents and others are about 8 percentage points more likely to express strong interest. These findings align with Mamkhezri [42] and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible [53]. They also reflect the stark policy differences between the two major political parties, particularly during the tenures of Presidents Biden and Trump. During his first term (2017–2021), Trump rolled back Obama-era fuel efficiency standards, slowing regulatory pressure on automakers to transition toward EVs. In his second term, the Trump administration has again taken a deregulatory approach. In contrast, the Biden administration adopted an aggressive pro-EV stance, prioritizing stricter fuel efficiency and emissions standards and allocating \$7.5 billion through the Infrastructure Investment and Jobs Act to develop a nationwide network of 500,000 EV charging stations.
@@
Income has a positive effect on EV purchase intent. Conditional on other factors, an additional \$50,000 in income changes the probability of strong interest (no interest) by 2.1–2.9 (2.2–3.6) percentage points. Our results find support in Zhao et al. [58] and Mamkhezri [42], who report greater support for EVs among higher-income individuals. The small covariate effect may reflect the growing availability of EVs across various market segments and the subsidies for low-income consumers, which reduce affordability barriers compared to earlier market years.